OpenStudy (anonymous):

Solve this equation for r: (2+3)r−6=31 A. r = 7.4 B. r = -31 C. r = 5 D. r = 11

OpenStudy (anonymous):

so 2 + 3 = 6 6r-6=31 31+6=37 6r=37 r=37/6

OpenStudy (anonymous):

my answer isn't there though

OpenStudy (anonymous):

if i were you i would just process of eilmination by putting the answers as r in the equation. For example: (2+3)(7.4)-6= ?

OpenStudy (anonymous):

and see if it equals 31